# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ 3.2 Equivalent characterization of order boundedness
The following lemma gives an equivalent characterization of the stochastic order by the comparison of cumulative distribution functions.
@@
($\Leftarrow$) By Hiai-Lawson-Lim [40, Proposition 3.11], it suffices to show $\int f\mathop{}\!\mathrm{d}\mu_{1}\leq\int f\mathop{}\!\mathrm{d}\mu_{2}$ for all bounded continuous increasing functions $f$. Let $f:\mathbb{R}\to\mathbb{R}$ be such a function, and we may assume $f$ is non-negative. Set
@@
and we have $f_{n}\uparrow f$ as $n\to\infty$. Note that the set $\{f>k/2^{n}\}$ is an open interval of the form $(x_{k},\infty)$, where (,):=assign($\infty$,$\infty$):$\emptyset$( ∞ , ∞ ) := ∅, since $f$ is increasing continuous. So we know that
@@
It remains to show that $\nu\geq_{\mathrm{st}}M$. Since $F_{\nu}=F=\inf_{\mu\in M}F_{\mu}$, we know that $F_{\nu}\leq F_{\mu}$ for all $\mu\in M$. Lemma 3.1 yields that $\nu\geq_{\mathrm{st}}\mu$ for all $\mu\in M$, which completes the proof. ∎
@@
If $M$ is an order bounded subset of $\mathcal{P}_{2}(\mathbb{R})$, then $M$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. In particular, any order interval $[\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$.
@@
$\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x),\ \text{ and }\ \int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x).$
@@
In particular, an order interval $[\mu,\nu]$ is closed due to the closedness of the stochastic order, so it is compact in $\mathcal{P}_{2}(\mathbb{R})$. ∎
@@
Since $\mathcal{W}_{2}(\mu_{n},\delta_{0})=(2/\log n)^{1/2}\to 0$ as $n\to\infty$, the set $\{\mu_{n}\}_{n=3}^{\infty}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. Note that
$\inf_{n\geq 3}F_{\mu_{n}}(x)=\left\{\begin{aligned} &0,&&x<0,\\ &1-\frac{1}{3^{2}\log 3},&&0\leq x<3,\\ &1-\frac{1}{k^{2}\log k},&&k-1\leq x<k,\ k\geq 4.\end{aligned}\right.$
@@
$\int_{0}^{\infty}\bigg{(}\sup_{n\geq 3}\mu_{n}\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u=\int_{0}^{\infty}\big{[}1-\inf_{n\geq 3}F_{\mu_{n}}(u)\big{]}u\mathop{}\!\mathrm{d}u\geq\sum_{k=3}^{\infty}\frac{k}{(k+1)^{2}\log(k+1)}=\infty,$
@@
Let $[\mu,\nu]$ be an order interval and $\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, $\rho\in[\mu,\nu]$. Then the following convergences are equivalent.
@@
It is obvious that (i)$\Rightarrow$(ii)$\Rightarrow$(iii), and we only need to show (iii)$\Rightarrow$(i).
By Proposition 3.4, we know that $[\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$. Then for any subsequence $\{\rho_{n_{k}}\}_{k\geq 1}\subset\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, there exists a subsequence $\{\rho_{n_{k_{j}}}\}_{j\geq 1}$ of $\{\rho_{n_{k}}\}_{k\geq 1}$ and a $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R})$ such that $\mathcal{W}_{2}(\rho_{n_{k_{j}}},\rho_{0})\to 0$ as $j\to\infty$. Thus, $\rho_{n_{k_{j}}}\to\rho_{0}$ weakly as $j\to\infty$. By (iii) and the uniqueness of the weak limit, we conclude $\rho_{0}=\rho$. Therefore, $\mathcal{W}_{2}(\rho_{n},\rho)\to 0$ as $n\to\infty$. ∎
@@ 4 Global Convergence to Order Interval
This section aims to prove Theorem 2.1 and Corollary 2.2. In Section 4.1, we target on the structure of global attractors of general semigroups on $\mathcal{P}_{2}(\mathbb{R})$ that preserve the stochastic order (namely, order-preserving semigroups). In Section 4.2, we prove a comparison theorem for extrinsic-law-dependent SDEs, and obtain the generation of order-preserving semigroup by the equation (1.1). The coming down from infinity property of hyper-dissipative McKean-Vlasov SDEs is shown in Section 4.3. At last, the proof of Theorem 2.1 and Corollary 2.2 is given in Section 4.4.
@@ 4.1 Global attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
In this subsection, we utilize the concepts of dissipative dynamical systems (see e.g., Hale [39]) to study the asymptotic behaviour of a semigroup on $\mathcal{P}_{2}(\mathbb{R})$ that preserves the stochastic order. With a Lyapunov condition, [14, 34] study the unique ergodicity of order-preserving Markov processes. Here, we focus on the structure of the global attractor of an order-preserving semigroup. Based on the existence of a globally attracting order bounded set, we prove that a global attractor of an order-preserving semigroup must be encompassed by two order related invariant measures, and these two invariant measures attract all solutions initiated from the regions above and below them with respect to the stochastic order, respectively (see Theorem 4.2).
@@
(i) is a direct consequence of the definition of $\omega$-limit sets. (ii)-(iv) are standard results for dissipative systems on complete metric spaces, for their proof one can see Hale [39, Lemma 3.1.1, Lemma 3.2.1 and Theorem 3.4.2 (ii)] respectively. ∎
Clearly, Lemma 4.1 (iii) implies that, if the global attractor $A$ exists, then $\omega(B)\subset A$ for any bounded set $B\subset\mathcal{P}_{2}(\mathbb{R})$. In the following, we give the definition of order-preserving semigroups.
@@
(iv). We only give the proof of the case of $\overline{\nu}$ (the case of $\underline{\nu}$ is similar). In fact, if $\rho\geq_{\mathrm{st}}\mu$ for some $\mu\in V$, by Lemma 4.1 (iii), it suffice to prove that $\omega(\rho)=\{\overline{\nu}\}$. Let $\pi\in\omega(\rho)$, Lemma 4.1 (i) entails that there exists a sequence $t_{k}\to\infty$ such that $\mathcal{W}_{2}(P^{*}_{t_{k}}\rho,\pi)\to 0$, as $k\to\infty$. Since $P^{*}$ is order-preserving and $\rho\geq_{\mathrm{st}}\mu$, we have $P^{*}_{t_{k}}\rho\geq_{\mathrm{st}}P^{*}_{t_{k}}\mu$ for any $k\geq 1$. Together with the closedness of the stochastic order, $\mathcal{W}_{2}(P^{*}_{t_{k}}\mu,\overline{\nu})\to 0$ as $k\to\infty$ implies that $\pi\geq_{\mathrm{st}}\overline{\nu}$. Since the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$, one has $\pi\in\omega(\rho)\subset\omega(K)\subset[\underline{\nu},\overline{\nu}]$. Therefore, $\pi=\overline{\nu}$. This proves (iv).
@@
(i). If we consider a semiflow preserving the stochastic order acting only on a closed subspace $B$ of $\mathcal{P}_{2}(\mathbb{R})$ (rather than the entire space $\mathcal{P}_{2}(\mathbb{R})$), by repeating the same proof, Theorem 4.2 still holds, where the order interval within $B$ is the restriction of the order interval to $B$. In fact, for semiflow preserving some closed partial order relation on a general complete metric space, by assuming the atttracting set $K$ is compact and order bounded, Theorem 4.2 still holds.
(ii). For semiflows preserving some closed partial order relation on a strongly ordered space, Hirsch [41, Theorem 3.3] implies the unique equilibrium must be globally attracting (Theorem 4.2 (vi)). However, Proposition A.1 shows that the strong ordering relation generated by $\leq_{\mathrm{st}}$ is empty. So, [41, Theorem 3.3] is not applicable here.
4.2 Comparison principle for extrinsic-law-dependent SDEs
@@
Next, we present a comparison theorem for extrinsic-law-dependent SDEs with respect to the stochastic order. Let $C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ be the space of continuous functions from $[s,\infty)$ to $\mathcal{P}_{2}(\mathbb{R})$. For any $\mu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $\xi\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$, we denote $\{X_{t}^{s,\mu,\xi}\}_{t\geq s}$ by the solution of the following extrinsic-law-dependent SDE
Then we have the following comparison principle.
Theorem 4.3.
Proof.
Since $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, Strassen’s theorem (see Lindvall [47, equation (3)]) yields that there exist $\tilde{\xi},\tilde{\eta}\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ such that
[…]
It suffices to show that $\mathcal{L}(X_{t}^{s,\mu,\tilde{\xi}})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{s,\nu,\tilde{\eta}})$ for all $t\in[s,s+T]$.
Note that
Since $\theta>0$, according to the comparison theorem for usual SDEs [37, Theorem 1.2], we know that
[…]
The desired result is proved. ∎
Remark 4.2.
Theorem 4.3 is not only used in the proof of the order-preservation for semigroups of McKean-Vlasov SDEs, but also crucial in the proof of our main result Theorem 2.3.
Associated to the McKean-Vlasov SDE (1.1), there is a semigroup $P_{t}^{*}$ acting on $\mathcal{P}_{2}(\mathbb{R})$ by
$P_{t}^{*}\mu:=\mathcal{L}(X_{t}^{\xi})\ \text{ with }\ \mathcal{L}(\xi)=\mu.$ (4.5)
In the following, we show that $P_{t}^{*}$ is an order-preserving semiflow.
Theorem 4.4.
Under Assumption 1, the semigroup $P^{*}:[0,\infty)\times\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ is an order-preserving semiflow.
Proof.
We first prove $P^{*}$ is order-preserving. It suffices to show that $P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho_{2}$ for any $\rho_{1}<_{\mathrm{st}}\rho_{2}$ and all $t\geq 0$. By Strassen’s theorem, there exist $\xi_{1},\xi_{2}\in L^{2}(\Omega,\mathcal{F}_{0},\mathbf{P})$ such that
$\mathcal{L}(\xi_{1})=\rho_{1},\ \ \mathcal{L}(\xi_{2})=\rho_{2},\ \ \text{ and}\ \ \xi_{1}\leq\xi_{2},\ \mathbf{P}-a.s.$
Since $\rho_{1}\neq\rho_{2}$, we have $\mathbf{P}\{\xi_{1}<\xi_{2}\}>0$ and hence
$\mathbf{E}[\xi_{1}]<\mathbf{E}[\xi_{2}].$ (4.6)
Let
$T:=\sup\{t>0:\mathbf{E}[X_{s}^{\xi_{1}}]<\mathbf{E}[X_{s}^{\xi_{2}}],\ \text{ for all }s\in[0,t]\}.$
Note that $\mathbf{E}[X_{t}^{\xi_{i}}]$ is continuous in $t$ for $i=1,2$, then it follows that $T>0$. The proof is divided into the following two cases:
$\textbf{Case 1}.\ \ T=\infty,\ \ \textbf{Case 2}.\ \ T<\infty.$
In case 1, let $\mu^{i}_{t}:=\mathcal{L}(X_{t}^{\xi_{i}}),t\in[0,\infty)$, $i=1,2$. Then we know that
$X_{t}^{\mu^{i},\xi_{i}}=X_{t}^{\xi_{i}},\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{t}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{t}(x),\ \text{ for all }\ t\geq 0.$
Then we conclude from Theorem 4.3 that
$P^{*}_{t}\rho_{1}=\mathcal{L}(X_{t}^{\mu^{1},\xi_{1}})\leq_{\mathrm{st}}\mathcal{L}(X_{t}^{\mu^{2},\xi_{2}})=P^{*}_{t}\rho_{2},\ \text{ for all }t\geq 0.$
In case 2, we have $0<T<\infty$. Let $\mu^{i}$ be defined as in case 1, we have $X_{t}^{\mu^{i},\xi_{i}}=X_{t}^{\xi_{i}}$ for all $t\geq 0$ and
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{t}(x)<\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{t}(x),\ t\in[0,T)\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{T}(x)=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{T}(x).$
Using Theorem 4.3 again, we deduce
$P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho_{2},\ \text{ for all }t\in[0,T].$
Note that $\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{T}(x)=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{T}(x)$, by the proof of (4.6), we conclude that $P^{*}_{T}\rho_{1}=P^{*}_{T}\rho_{2}$. Then we have $P^{*}_{t}\rho_{1}=P^{*}_{t}\rho_{2}$ for all $t\geq T$.
So the order-preservation of $P^{*}$ is proved. It remains to show that $P^{*}$ is jointly continuous, which follows from [48, Proposition 2.3] directly. ∎
4.3 Coming down from infinity
@@
Indeed, we want to show the solutions are $p$-integrable for all $p\geq 2$ if the initial conditions are just square integrable. For this purpose, we need the following extrinsic-law case.
@@
$\begin{split}&\textbf{Case 1:}\ y_{t}\geq\frac{\alpha\delta}{2}t,\ \text{ if }\ y_{s}\leq A_{t},\ \text{ for all }\ s\in[0,t];\\ &\textbf{Case 2:}\ y_{t}\geq A_{t},\ \text{ if there exists $s\in[0,t]$ such that }\ y_{s}\geq A_{t}.\end{split}$
@@
Denote $\mu_{t}:=\mathcal{L}(X_{t}^{\xi})$ for $t\geq 0$, and by Lemma 4.5 and Lemma 4.6, we know that for any $t>0$,
@@
where the constant $C(p)$ come from Lemma 4.7. Clearly, $\delta_{0}\in K$, so $K$ is nonempty. We now show (i)-connectedness. Because $K$ is convex and the operation of convex combinations is continuous in $\mathcal{P}_{2}(\mathbb{R})$, we obtain $K$ is connected. Then we want to show (iii). Since $K$ is bounded in $\mathcal{P}_{p}(\mathbb{R})$ for all $p>2$, it follows from Corollary 3.3 that $K$ is order bounded. By Proposition 3.4, to prove (i)-compactness, we only need to show $K$ is closed. Suppose a $\mathcal{P}_{2}$-Cauchy sequence $\{\mu_{n}\}_{n\geq 1}\subset K$ converges to a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$, and it remains to check $\mu\in K$, i.e., $\mu\in K_{p}$ for all $p\geq 2$. By [68, Theorem 6.9], $\mu_{n}$ weakly converges to $\mu$. For any $p\geq 2$, since $\{\mu_{n}\}_{n\geq 1}\subset K_{p}$, then
@@
We first give the following lemma, which is a direct consequence of Feng-Qu-Zhao [33, Theorem 4.8]. Let $C_{b}(\mathbb{R};\mathcal{P}_{2}(\mathbb{R}))$ be the collection of all bounded continuous paths $\{\nu_{t}\}_{t\in\mathbb{R}}$ in $\mathcal{P}_{2}(\mathbb{R})$.
@@
Obviously, $\mu_{i}$ is an invariant measures of (4.16). Denote by $X_{t}^{i,x}$ the solution of (4.16) with initial value $x$ and starting time 00, $i=1,2$. Since $m_{1}\leq m_{2}$, it follows from the comparison theorem (Theorem 4.3) that
@@
By Lemma 4.9 (ii), $\mu_{i}$ is the unique invariant measure of (4.16) and
@@
Next, we want to show the convergence in $\mathcal{P}_{2}(\mathbb{R})$. Observe that, for any bounded continuous function $f:\mathbb{R}\to\mathbb{R}$, with $|f|_{\infty}:=\sup_{x\in\mathbb{R}}|f(x)|$,
$\Big{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{2}(x)\Big{|}\leq|f|_{\infty}d_{2}\big{(}\nu_{1},\nu_{2}\big{)},$ (4.19)
and
$\Big{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{2}(x)\Big{|}\leq d_{2}\big{(}\nu_{1},\nu_{2}\big{)}.$ (4.20)
Combine (4.18)-(4.20), and it follows from the equivalent characterization of $\mathcal{W}_{2}$-convergence (see Villani [68, Theorem 6.9]) that

### v2

@@ 2.2 Equivalent characterization of order boundedness
We begin with an equivalent characterization of the stochastic order in $\mathcal{P}_{2}(\mathbb{R})$ by comparison of cumulative distribution functions.
@@
($\Leftarrow$) By Hiai-Lawson-Lim [25, Proposition 3.11], it suffices to show $\int f\mathop{}\!\mathrm{d}\mu_{1}\leq\int f\mathop{}\!\mathrm{d}\mu_{2}$ for all bounded continuous increasing functions $f$. Let $f:\mathbb{R}\to\mathbb{R}$ be such a function, and we may assume $f$ is non-negative. Set
@@
and we have $f_{n}\uparrow f$ as $n\to\infty$. Note that the set $\{f>k/2^{n}\}$ is an open interval of the form $(x_{k},\infty)$, where $(\infty,\infty):=\emptyset$, since $f$ is increasing continuous. So we have
@@
It remains to show that $\nu\geq_{\mathrm{st}}M$. Since $F_{\nu}=F=\inf_{\mu\in M}F_{\mu}$, we have $F_{\nu}\leq F_{\mu}$ for all $\mu\in M$. Lemma 2.1 yields that $\nu\geq_{\mathrm{st}}\mu$ for all $\mu\in M$, which completes the proof. ∎
@@
Remark 2.3.
We remark here that in higher dimensional case, there is also an equivalent characterization of order boundedness for a subset of $\mathcal{P}_{2}(\mathbb{R}^{d})$:
A subset $M\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ is order bounded from above (below) in $\mathcal{P}_{2}(\mathbb{R}^{d})$ if and only if the marginals of $M$ on each component are order bounded from above (below).
We give a brief proof here. ($\Rightarrow$) is obvious, and we only show ($\Leftarrow$). We will find a lower bound $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ of $M$, and the construction of an upper bound is similar, which is omitted.
[…]
$\hat{\mu}\big((-\infty,u]\big)=\mu\big(\mathbb{R}^{d}\setminus(u,\infty)^{d}\big)\leq\sum_{i=1}^{d}\mu\big(\mathbb{R}^{i-1}\times(-\infty,u]\times\mathbb{R}^{d-i}\big).$ (2.11)
Given that marginals of $M$ on each component are order bounded below, Proposition 2.2 (ii) implies, for each $i=1,2,\dots,d$,
$\int_{0}^{\infty}\bigg(\sup_{\mu\in M}\mu\big(\mathbb{R}^{i-1}\times(-\infty,-u]\times\mathbb{R}^{d-i}\big)\bigg)u\mathop{}\!\mathrm{d}u<\infty.$ (2.12)
Combining (2.11)(2.12) yields
$\int_{0}^{\infty}\bigg(\sup_{\mu\in M}\hat{\mu}\big((-\infty,-u]\big)\bigg)u\mathop{}\!\mathrm{d}u\leq\sum_{i=1}^{d}\int_{0}^{\infty}\bigg(\sup_{\mu\in M}\mu\big(\mathbb{R}^{i-1}\times(-\infty,-u]\times\mathbb{R}^{d-i}\big)\bigg)u\mathop{}\!\mathrm{d}u<\infty,$
so there is some $\nu\in\mathcal{P}_{2}(\Delta)$ such that $\nu\leq_{\mathrm{st}}\hat{\mu}$ for all $\mu\in M$. Now we observe $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and $\nu\leq_{\mathrm{st}}\hat{\mu}\leq_{\mathrm{st}}\mu$ for all $\mu\in M$. This finishes the proof.
@@
If $M$ is an order bounded subset of $\mathcal{P}_{2}(\mathbb{R})$, then $M$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. In particular, any order interval $\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$.
@@
[…]
@@
In particular, an order interval $\mu,\nu]$ is closed due to the closedness of the stochastic order, so it is compact in $\mathcal{P}_{2}(\mathbb{R})$. ∎
@@
Since $\mathcal{W}_{2}(\mu_{n},\delta_{0})=(2/\log n)^{1/2}\to 0$ as $n\to\infty$, the set $\{\mu_{n}\}_{n=2}^{\infty}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. Note that
$\inf_{n\geq 2}F_{\mu_{n}}(x)=\left\{\begin{aligned} &0,&&x<0,\\ &1-\frac{1}{2^{2}\log 2},&&0\leq x<2,\\ &1-\frac{1}{k^{2}\log k},&&k-1\leq x<k,\ k\geq 3.\end{aligned}\right$.
@@
$\int_{0}^{\infty}\bigg(\sup_{n\geq 2}\mu_{n}\big((u,\infty)\big)\bigg)u\mathop{}\!\mathrm{d}u=\int_{0}^{\infty}\big[1-\inf_{n\geq 2}F_{\mu_{n}}(u)\big]u\mathop{}\!\mathrm{d}u\geq\sum_{k=2}^{\infty}\frac{k}{(k+1)^{2}\log(k+1)}=\infty,$
@@
Let $\mu,\nu]$ be an order interval and $\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, $\rho\in[\mu,\nu]$. Then the following convergences are equivalent.
@@
[…]
By Proposition 2.4, we have $\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$. Then for any subsequence $\{\rho_{n_{k}}\}_{k\geq 1}\subset\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, there exists a subsequence $\{\rho_{n_{k_{j}}}\}_{j\geq 1}$ of $\{\rho_{n_{k}}\}_{k\geq 1}$ and a $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ such that $\mathcal{W}_{2}(\rho_{n_{k_{j}}},\rho_{0})\to 0$ as $j\to\infty$. Thus, $\rho_{n_{k_{j}}}\to\rho_{0}$ weakly as $j\to\infty$. By (iii) and the uniqueness of the weak limit, we conclude $\rho_{0}=\rho$. Therefore, $\mathcal{W}_{2}(\rho_{n},\rho)\to 0$ as $n\to\infty$. ∎
@@ 3 Global Convergence to Order Interval
This section aims to prove Theorem 1.2 and Corollary 1.3. In Section 3.1, we target on the structure of global attractors of general semigroups on $\mathcal{P}_{2}(\mathbb{R})$ that preserve the stochastic order (namely, order-preserving semigroups). The coming down from infinity property of hyper-dissipative McKean-Vlasov SDEs is shown in Section 3.2. At last, the proof of Theorem 1.2 and Corollary 1.3 is given in Section 3.3.
@@ 3.1 Global attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
The notion of order-preserving has already appeared in the study of uniqueness of invariant measures, ergodicity, and synchronization phenomena for certain classes of Markov processes; see, for instance, [5, 37, 46]. In this subsection, we instead utilize concepts from the theory of dissipative dynamical systems (see, e.g., Hale [24]) to study the asymptotic behaviour of a general semigroup on $\mathcal{P}_{2}(\mathbb{R})$ that preserves the stochastic order, allowing in particular for the presence of multiple invariant measures. Based on the existence of a globally attracting order bounded set, we prove that a global attractor of an order-preserving semigroup must be encompassed by two order related invariant measures, and these two invariant measures attract all solutions initiated from the regions above and below them with respect to the stochastic order, respectively (see Proposition 3.2).
@@
(i) is a direct consequence of the definition of $\omega$-limit sets. (ii)-(iv) are standard results for dissipative systems on complete metric spaces, for their proof one can see Hale [24, Lemma 3.1.1, Lemma 3.2.1 and Theorem 3.4.2 (ii)] respectively. ∎
Clearly, Lemma 3.1 (iii) implies that, if the global attractor $A$ exists, then $\omega(B)\subset A$ for any bounded set $B\subset\mathcal{P}_{2}(\mathbb{R})$. In the following, we give the definition of order-preserving semiflows.
@@
(iv). We only give the proof of the case of $\overline{\nu}$ (the case of $\underline{\nu}$ is similar). In fact, if $\rho\geq_{\mathrm{st}}\mu$ for some $\mu\in V$, by Lemma 3.1 (iii), it suffice to prove that $\omega(\rho)=\{\overline{\nu}\}$. Let $\pi\in\omega(\rho)$, Lemma 3.1 (i) entails that there exists a sequence $t_{k}\to\infty$ such that $\mathcal{W}_{2}(P^{*}_{t_{k}}\rho,\pi)\to 0$, as $k\to\infty$. Since $P^{*}$ is order-preserving and $\rho\geq_{\mathrm{st}}\mu$, we have $P^{*}_{t_{k}}\rho\geq_{\mathrm{st}}P^{*}_{t_{k}}\mu$ for any $k\geq 1$. Together with the closedness of the stochastic order, $\mathcal{W}_{2}(P^{*}_{t_{k}}\mu,\overline{\nu})\to 0$ as $k\to\infty$, one has $\pi\geq_{\mathrm{st}}\overline{\nu}$. Since the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$, one has $\pi\in\omega(\rho)\subset\omega(K)\subset[\underline{\nu},\overline{\nu}]$. Therefore, $\pi=\overline{\nu}$. This proves (iv).
@@
(i). If we consider a semiflow preserving the stochastic order acting only on a closed subspace $B$ of $\mathcal{P}_{2}(\mathbb{R})$ (rather than the entire space $\mathcal{P}_{2}(\mathbb{R})$), by repeating the same proof, Proposition 3.2 still holds, where the order interval within $B$ is the restriction of the order interval to $B$. In fact, for semiflow preserving some closed partial order relation on a general complete metric space, by assuming the atttracting set $K$ is compact and order bounded, Proposition 3.2 still holds.
(ii). For semiflows preserving some closed partial order relation on a strongly ordered space, Hirsch [27, Theorem 3.3] implies the unique equilibrium must be globally attracting (Proposition 3.2 (vi)). However, Proposition A.1 shows that the strong ordering relation generated by $\leq_{\mathrm{st}}$ is empty. So, [27, Theorem 3.3] is not applicable here.
3.2 Coming down from infinity
@@
Indeed, we want to show the solutions are $p$-integrable for all $p\geq 2$ if the initial conditions are just square integrable. For this purpose, we need the following extrinsic-law case. Let $C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ be the space of continuous functions from $s,\infty)$ to $\mathcal{P}_{2}(\mathbb{R})$. For any $\mu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $\xi\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$, we denote $\{X_{t}^{s,\mu,\xi}\}_{t\geq s}$ by the solution of the following extrinsic-law-dependent SDE
@@
[…]
@@
Denote $\mu_{t}:=\mathcal{L}(X_{t}^{\xi})$ for $t\geq 0$, and by Lemma 3.3 and Lemma 3.4, we have for any $t>0$,
@@
where the constant $C(p)$ comes from Lemma 3.5. Clearly, $\delta_{0}\in K$, so $K$ is nonempty. We now show (i)-connectedness. Because $K$ is convex and the operation of convex combinations is continuous in $\mathcal{P}_{2}(\mathbb{R})$, we obtain $K$ is connected. Then we want to show (iii). Since $K$ is bounded in $\mathcal{P}_{p}(\mathbb{R})$ for all $p>2$, it follows from Corollary 2.3 that $K$ is order bounded. By Proposition 2.4, to prove (i)-compactness, we only need to show $K$ is closed. Suppose a $\mathcal{P}_{2}$-Cauchy sequence $\{\mu_{n}\}_{n\geq 1}\subset K$ converges to a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$, and it remains to check $\mu\in K$, i.e., $\mu\in K_{p}$ for all $p\geq 2$. By [54, Theorem 6.9], $\mu_{n}$ weakly converges to $\mu$. For any $p\geq 2$, since $\{\mu_{n}\}_{n\geq 1}\subset K_{p}$, then
@@
We first give the following lemma, which is a direct consequence of Feng-Qu-Zhao [17, Theorem 4.8]. Let $C_{b}(\mathbb{R};\mathcal{P}_{2}(\mathbb{R}))$ be the collection of all bounded continuous paths $\{\nu_{t}\}_{t\in\mathbb{R}}$ in $\mathcal{P}_{2}(\mathbb{R})$.
@@
Obviously, $\mu_{i}$ is an invariant measures of (3.11). Denote by $X_{t}^{i,x}$ the solution of (3.11) with initial value $x$ and starting time 0, $i=1,2$. Since $m_{1}\leq m_{2}$, it follows from the comparison theorem for usual SDEs [20, Theorem 1.2] that
@@
By Lemma 3.7, $\mu_{i}$ is the unique invariant measure of (3.11) and
@@
Note that $\mathcal{W}_{2}(\mu,\nu)\leq\sqrt{2d_{2}(\mu,\nu)}$ (see e.g. [54, Theorem 6.15]). Then we conclude that